# Two-dimensional FrBD friction models for rolling contact

### v1

@@ 3.4 Linear rolling contact model for large spin slips
$\tilde{\mathbf{\Sigma}}_{\varphi}(\bm{x},s)$ $\triangleq\mathbf{\Sigma}\bigl(\bar{\bm{v}}(\bm{x},s),s\bigr)+\tilde{\mathbf{H}}(\bm{x},s)\mathbf{A}_{\varphi_{2}}(s).$ (54c)
@@ 4.2 Dissipativity and passivity
$\begin{split}\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&=\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=\iint_{\mathscr{C}}p(\bm{x})\bm{f}^{\mathrm{T}}(\bm{x},s)\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &=\iint_{\mathscr{C}}p(\bm{x})\bm{f}^{\mathrm{T}}(\bm{x},s)\bm{\sigma}(s)\operatorname{d\!}\bm{x}+\iint_{\mathscr{C}}p(\bm{x})\bigl[xf_{y}(\bm{x},s)-yf_{x}(\bm{x},s)\bigr]\varphi(s)\operatorname{d\!}\bm{x}\\ &=\bm{F}_{\bm{x}}^{\mathrm{T}}(s)\bm{\sigma}(s)+M_{z}(s)\varphi(s),\qquad s\in[0,S].\end{split}$ (64)
@@
$\begin{split}\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&=\bm{F}_{\bm{x}}^{\mathrm{T}}(s)\bm{\sigma}(s)+M_{z}(s)\varphi(s)\\ &\qquad-\varphi(s)\iint_{\mathscr{C}}p(\bm{x})\bigl[z_{x}(\bm{x},s)f_{y}(\bm{x},s)-z_{y}(\bm{x},s)f_{x}(\bm{x},s)\bigr]\operatorname{d\!}\bm{x},\qquad s\in[0,S],\end{split}$ (65)
@@
$\begin{split}&\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\bm{z}(\cdot,s),\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=\bm{F}_{\bm{x}}^{\mathrm{T}}(s)\bm{\sigma}(s)+M_{z}(s)\varphi(s)\\ &\hskip 17.00024pt\hskip 17.00024pt\qquad+\varphi_{\psi}(s)\iint_{\mathscr{C}}p(\bm{x})\bigl[z_{x}(\bm{x},s)f_{y}(\bm{x},s)-z_{y}(\bm{x},s)f_{x}(\bm{x},s)\bigr]\operatorname{d\!}\bm{x},\qquad s\in[0,S].\end{split}$ (66)
@@
$\begin{split}&\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\bm{z}(\cdot,s),\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=\bm{F}_{\bm{x}}^{\mathrm{T}}(s)\bm{\sigma}(s)+M_{z}(s)\varphi(s)\\ &\hskip 17.00024pt\hskip 17.00024pt\qquad-\varphi_{\gamma}(s)\iint_{\mathscr{C}}p(\bm{x})\bigl[z_{x}(\bm{x},s)f_{y}(\bm{x},s)-z_{y}(\bm{x},s)f_{x}(\bm{x},s)\bigr]\operatorname{d\!}\bm{x},\qquad s\in[0,S].\end{split}$ (67)
Equations (66) and (67), which apply when one of the bodies in rolling contact is rigid, differ in how they treat the geometric and effective spin terms. In essence, Eq. (66) includes an additional contribution from the turn spin $\varphi_{\psi}(s)$, arising from the fact that the bristle tips experience a relative velocity component associated with an effective rotation about the vertical axis. In contrast, Eq. (67) states that the geometric spin $\varphi_{\gamma}(s)$ does not dissipate power in the undeformed configuration, since it affects only the trajectory of the bristles within the contact patch and is therefore related to the velocity of their roots rather than their tips (which only feel the effect of $\varphi_{\psi}(s)$). Both Eq. (66) and (67) are simultaneously valid; the difference lies solely in the manner in which the vertical moment is evaluated. It is also worth remarking that, for Model 3.2, the equivalence $\bar{\bm{v}}_{\textnormal{r}}(\bm{z}(\bm{x},s),\bm{x},s)=\bar{\bm{v}}(\bm{x},s)$ only holds when $\varphi_{2}(s)=0$. Since, from a physical viewpoint, passivity should be regarded as a property of the friction model in isolation, it appears reasonable to consider the case $\varphi_{2}(s)=0$ when analysing Model 3.2. Indeed, when $\varphi_{2}(s)\not=0$, Model 3.2 may be interpreted as a feedback system where the rigid relative velocity depends upon the state through a feedback term $\mathbf{A}_{\varphi_{2}}(s)\bm{z}(\bm{x},s)$. In this context, a question that naturally arises is whether all the nondimensional velocities $\bar{\bm{v}}(\bm{x},s)$, rolling speeds $V_{\textnormal{r}}(s)$, and spins $\bm{\varphi}(s)$ (with $\varphi_{2}(s)=0$ in Eq. (57)) ensuring the well-posedness of the PDEs (45) and (57), as asserted by Theorems 3.1 and 3.2, imply automatically $\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}\geq 0$ for all $s\in[0,S]$ in (64). In particular, limited to classical solutions of the equations under consideration, the concepts of dissipativity and passivity are enounced according to Definition 4.3 below. A more general introduction to these notions, limited to linear systems, is offered in [82].
@@
$\begin{split}&\int_{0}^{s}w\bigl(\bm{f}(\cdot,s^{\prime}),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s^{\prime})\bigr)\operatorname{d\!}s^{\prime}=\int_{0}^{s}w\bigl(\bm{f}(\cdot,s^{\prime}),\bar{\bm{v}}(\cdot,s^{\prime})\bigr)\operatorname{d\!}s^{\prime}\geq W\bigl(\bm{z}(\cdot,s)\bigr)-W\bigl(\bm{z}_{0}(\cdot)\bigr),\qquad s\in[0,S].\end{split}$ (68)
It is called passive if $w(\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s))=\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}$.
@@ 5.1.1 Force-slip surfaces
where $(\hat{\bm{F}}_{\bm{x}},\hat{M}_{z})\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{4}\mapsto\mathbb{R}^{3}$ may be generally referred to as the force-slip surfaces. Indeed, for a given value of the rolling speed $V_{\textnormal{r}}$, regarding the slip and spin inputs $(\bm{\sigma},\bm{\varphi})$ as independent variables, each steady-state characteristic, namely the longitudinal force $F_{x}$, the lateral force $F_{y}$, and the vertical moment $M_{z}$, may be interpreted as a four-dimensional surface living in a five-dimensional space. Typically, the force-slip surfaces are usually continuous, uniformly bounded, and at least locally $C^{1}$. In this context, it is worth observing that the rolling speed itself could be regarded as an input; in which case, the force-slip surfaces should be more properly interpreted as five-dimensional surfaces embedded in a six-dimensional space. However, most friction and rolling contact models assume a minor dependence upon $V_{\textnormal{r}}$ compared to the slips and spins.
@@ 6 Conclusions
Beyond the specific formulations presented in the paper, the developed methodology, combining physical-oriented reasoning, implicit-function-based inversion, and distributed PDE representations, establishes a general template for future extensions. Potential directions include incorporating more detailed bristle mechanics, fully nonlinear damping laws, or coupling with flexible bodies. Indeed, the results of this work open the way for new physically-grounded, control-oriented rolling contact representations and provide a rigorous foundation for next-generation friction modelling in applications ranging from vehicle dynamics to tribology. Finally, starting with the formulations introduced in this manuscript, model-order-reduction strategies could be explored for potential integration with larger, system-level dynamic simulations of complex mechanical systems.
@@
L.R. is the sole author of and contributor to the manuscript.
@@
$\begin{split}\dot{W}(s)&\leq-\dfrac{1}{2}\iint_{\mathscr{C}}(\theta\eta-2M_{\tilde{\mathbf{\Sigma}}_{\varphi}}-1)P(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}+\dfrac{1}{2}M_{\mathbf{H}}^{2}\iint_{\mathscr{C}}P(\bm{x})\mathinner{\!\left\lVert\bm{v}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (90)
@@
$\begin{split}\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}+\iint_{\mathscr{C}}p(\bm{x})\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}^{\mathrm{T}}(\bm{x},s)}{\operatorname{d\!}{s}}}V_{\textnormal{r}}\mathbf{\Sigma}_{1}\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\Bigl[\mathbf{\Sigma}_{0}+V_{\textnormal{r}}\tilde{\mathbf{\Sigma}}(\bm{x},s)^{\mathrm{T}}\mathbf{\Sigma}_{1}\Bigr]\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad+\iint_{\mathscr{C}}V_{\textnormal{r}}p(\bm{x})\bar{\bm{v}}^{\mathrm{T}}(\bm{x},s)\mathbf{H}\bigl(\bar{\bm{v}}(\bm{x},s),s\bigr)^{\mathrm{T}}\mathbf{\Sigma}_{1}\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (93)
@@
$\begin{split}&\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}-\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{Q}_{1}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\hskip 17.00024pt+\iint_{\mathscr{C}}V_{\textnormal{r}}p(\bm{x})\bar{\bm{v}}^{\mathrm{T}}(\bm{x},s)\mathbf{Q}_{2}(\bm{x},s)\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\hskip 17.00024pt+\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S),\end{split}$ (95)
@@
$\mathbf{Q}_{2}(\bm{x},s)$ $\triangleq\mathbf{H}\bigl(\bar{\bm{v}}(\bm{x},s),s\bigr)^{\mathrm{T}}\mathbf{\Sigma}_{1}=\mathbf{M}^{2}\bigl(V_{\textnormal{r}}\bar{\bm{v}}(\bm{x},s)\bigr)\mathbf{G}^{-1}\bigl(V_{\textnormal{r}}\bar{\bm{v}}(\bm{x},s)\bigr)\mathbf{\Sigma}_{1}.$ (96b)
@@
$\begin{split}\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)+\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (97)
@@
$\begin{split}\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)-\dfrac{1}{2}\iint_{\mathscr{C}}\bigl(\nabla_{\bm{x}}\cdot p(\bm{x})\bar{\bm{V}}(\bm{x})\bigr)\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (98)
@@
$\begin{split}\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\dot{W}\bigl(\bm{z}(\cdot,s)\bigr),\qquad s\in(0,S).\end{split}$ (99)
Integrating the above Eq. (99) proves (68) with $w(\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s))=\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}$. ∎

### v2

@@ 3.4 Linear rolling contact model for large spin slips
$\tilde{\mathbf{\Sigma}}_{\varphi}(\bm{x},s)$ $\triangleq\mathbf{\Sigma}\bigl(\bar{\bm{v}}(\bm{x},s),s\bigr)-\tilde{\mathbf{H}}(\bm{x},s)\mathbf{A}_{\varphi_{2}}(s).$ (56c)
@@ 4.2 Dissipativity and passivity
$\begin{split}-\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&=-\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=-\iint_{\mathscr{C}}p(\bm{x})\bm{f}^{\mathrm{T}}(\bm{x},s)\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &=\iint_{\mathscr{C}}p(\bm{x})\bm{f}^{\mathrm{T}}(\bm{x},s)\bm{\sigma}(s)\operatorname{d\!}\bm{x}+\iint_{\mathscr{C}}p(\bm{x})\bigl[xf_{y}(\bm{x},s)-yf_{x}(\bm{x},s)\bigr]\varphi(s)\operatorname{d\!}\bm{x}\\ &=\bm{F}_{\bm{x}}^{\mathrm{T}}(s)\bm{\sigma}(s)+M_{z}(s)\varphi(s),\qquad s\in[0,S].\end{split}$ (66)
@@
$\begin{split}-\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&=\bm{F}_{\bm{x}}^{\mathrm{T}}(s)\bm{\sigma}(s)+M_{z}(s)\varphi(s)\\ &\qquad-\varphi(s)\iint_{\mathscr{C}}p(\bm{x})\bigl[z_{x}(\bm{x},s)f_{y}(\bm{x},s)-z_{y}(\bm{x},s)f_{x}(\bm{x},s)\bigr]\operatorname{d\!}\bm{x},\qquad s\in[0,S],\end{split}$ (67)
@@
$\begin{split}&-\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\bm{z}(\cdot,s),\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=\bm{F}_{\bm{x}}^{\mathrm{T}}(s)\bm{\sigma}(s)+M_{z}(s)\varphi(s)\\ &\hskip 17.00024pt\hskip 17.00024pt\qquad+\varphi_{\psi}(s)\iint_{\mathscr{C}}p(\bm{x})\bigl[z_{x}(\bm{x},s)f_{y}(\bm{x},s)-z_{y}(\bm{x},s)f_{x}(\bm{x},s)\bigr]\operatorname{d\!}\bm{x},\qquad s\in[0,S].\end{split}$ (68)
@@
$\begin{split}&-\bigl\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\bm{z}(\cdot,s),\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=\bm{F}_{\bm{x}}^{\mathrm{T}}(s)\bm{\sigma}(s)+M_{z}(s)\varphi(s)\\ &\hskip 17.00024pt\hskip 17.00024pt\qquad-\varphi_{\gamma}(s)\iint_{\mathscr{C}}p(\bm{x})\bigl[z_{x}(\bm{x},s)f_{y}(\bm{x},s)-z_{y}(\bm{x},s)f_{x}(\bm{x},s)\bigr]\operatorname{d\!}\bm{x},\qquad s\in[0,S].\end{split}$ (69)
Equations (68) and (69), which apply when one of the bodies in rolling contact is rigid, differ in how they treat the geometric and effective spin terms. In essence, Eq. (68) includes an additional contribution from the turn spin $\varphi_{\psi}(s)$, arising from the fact that the bristle tips experience a relative velocity component associated with an effective rotation about the vertical axis. In contrast, Eq. (69) states that the geometric spin $\varphi_{\gamma}(s)$ does not dissipate power in the undeformed configuration, since it affects only the trajectory of the bristles within the contact patch and is therefore related to the velocity of their roots rather than their tips (which only feel the effect of $\varphi_{\psi}(s)$). Both Eq. (68) and (69) are simultaneously valid; the difference lies solely in the manner in which the vertical moment is evaluated. It is also worth remarking that, for Model 3.2, the equivalence $\bar{\bm{v}}_{\textnormal{r}}(\bm{z}(\bm{x},s),\bm{x},s)=\bar{\bm{v}}(\bm{x},s)$ only holds when $\varphi_{2}(s)=0$. Since, from a physical viewpoint, passivity should be regarded as a property of the friction model in isolation, it appears reasonable to consider the case $\varphi_{2}(s)=0$ when analysing Model 3.2. Indeed, when $\varphi_{2}(s)\not=0$, Model 3.2 may be interpreted as a feedback system where the rigid relative velocity depends upon the state through a feedback term $\mathbf{A}_{\varphi_{2}}(s)\bm{z}(\bm{x},s)$. In this context, a question that naturally arises is whether all the nondimensional velocities $\bar{\bm{v}}(\bm{x},s)$, rolling speeds $V_{\textnormal{r}}(s)$, and spins $\bm{\varphi}(s)$ (with $\varphi_{2}(s)=0$ in Eq. (59)) ensuring the well-posedness of the PDEs (47) and (59), as asserted by Theorems 3.1 and 3.2, imply automatically $\langle p(\bm{x})\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}\geq 0$ for all $s\in[0,S]$ in (66). In particular, restricted to classical solutions of the equations under consideration, the concepts of dissipativity and passivity are enounced according to Definition 4.3 below. A more general introduction to these notions, limited to linear systems, is offered in [82].
@@
$\begin{split}&\int_{0}^{s}w\bigl(\bm{f}(\cdot,s^{\prime}),-\bar{\bm{v}}_{\textnormal{r}}(\cdot,s^{\prime})\bigr)\operatorname{d\!}s^{\prime}=\int_{0}^{s}w\bigl(\bm{f}(\cdot,s^{\prime}),-\bar{\bm{v}}(\cdot,s^{\prime})\bigr)\operatorname{d\!}s^{\prime}\geq W\bigl(\bm{z}(\cdot,s)\bigr)-W\bigl(\bm{z}_{0}(\cdot)\bigr),\qquad s\in[0,S].\end{split}$ (70)
It is called passive if $w(\bm{f}(\cdot,s),-\bar{\bm{v}}(\cdot,s))=-\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=-\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}$.
@@ 5.1.1 Force-slip surfaces
where $(\hat{\bm{F}}_{\bm{x}},\hat{M}_{z})\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{4}\mapsto\mathbb{R}^{3}$ may be generally referred to as the force-slip surfaces. Indeed, for a given value of the rolling speed $V_{\textnormal{r}}$, regarding the slip and spin inputs $(\bm{\sigma},\bm{\varphi})$ as independent variables, each steady-state characteristic, namely the longitudinal force $F_{x}$, the lateral force $F_{y}$, and the vertical moment $M_{z}$, may be interpreted as a four-dimensional surface living in a five-dimensional space [^7] The mapping $(\bm{\sigma},\bm{\varphi})\mapsto(\bm{F}_{\bm{x}},M_{z})$ is from $\mathbb{R}^{4}$ into $\mathbb{R}^{3}$. Thus, considering, for instance, $F_{x}\in\mathbb{R}$, $\mathrm{graph}(\bm{\sigma},\bm{\varphi},F_{x})\subset\mathbb{R}^{5}$ lives in a five-dimensional space. Regarding also $V_{\textnormal{r}}\in\mathbb{R}$ as a variable, $\mathrm{graph}(\bm{\sigma},\bm{\varphi},V_{\textnormal{r}},F_{x})\subset\mathbb{R}^{6}$ lives instead in a six-dimensional space. The situation is analogous for the other characteristics.. Typically, the force-slip surfaces are usually continuous, uniformly bounded, and at least locally $C^{1}$. In this context, it is worth observing that the rolling speed itself could be regarded as an input; in which case, the force-slip surfaces should be more properly interpreted as five-dimensional surfaces embedded in a six-dimensional space. However, most friction and rolling contact models assume a minor dependence upon $V_{\textnormal{r}}$ compared to the slips and spins.
@@ 6 Conclusions
Beyond the specific formulations presented in the paper, the developed methodology, combining physical-oriented reasoning, implicit-function-based inversion, and distributed PDE representations, establishes a general template for future extensions. Potential directions include incorporating more detailed bristle mechanics, fully nonlinear damping laws, or coupling with flexible bodies. Indeed, the results of this work open the way for new physically-grounded, control-oriented rolling contact representations and provide a rigorous foundation for next-generation friction modelling in applications ranging from vehicle dynamics to tribology. In particular, a limitation of the specific model implementation presented in this paper relates to the simplicity of the assumed rheological representation of the bristle element, which may be appropriate for elastic or viscoelastic solids where relaxation effects are minor, but cannot accurately capture complex dissipative behaviours observed, for instance, in polymers and composite materials. Finally, starting with the formulations introduced in this manuscript, model-order-reduction strategies could be explored for potential integration with larger, system-level dynamic simulations of complex mechanical systems.
@@
Acknowledgements
The author thanks both reviewers for their valuable suggestions and comments, with particular appreciation to Reviewer 2 for especially helpful feedback that substantially improved the manuscript.
@@ Acknowledgements
L.R. is the sole author and contributor to the manuscript.
@@
$\begin{split}\dot{W}(s)&\leq-\dfrac{1}{2}\iint_{\mathscr{C}}(\theta\eta-2M_{\tilde{\mathbf{\Sigma}}_{\varphi}}-1)P(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}+\dfrac{1}{2}M_{\mathbf{H}}^{2}\iint_{\mathscr{C}}P(\bm{x})\mathinner{\!\left\lVert\bar{\bm{v}}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (92)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&=-\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}-\iint_{\mathscr{C}}p(\bm{x})\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}^{\mathrm{T}}(\bm{x},s)}{\operatorname{d\!}{s}}}V_{\textnormal{r}}\mathbf{\Sigma}_{1}\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &=-\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\Bigl[\mathbf{\Sigma}_{0}+V_{\textnormal{r}}\tilde{\mathbf{\Sigma}}(\bm{x},s)^{\mathrm{T}}\mathbf{\Sigma}_{1}\Bigr]\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad-\iint_{\mathscr{C}}V_{\textnormal{r}}p(\bm{x})\bar{\bm{v}}^{\mathrm{T}}(\bm{x},s)\mathbf{H}\bigl(\bar{\bm{v}}(\bm{x},s),s\bigr)^{\mathrm{T}}\mathbf{\Sigma}_{1}\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (95)
@@
$\begin{split}&-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}-\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{Q}_{1}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\hskip 17.00024pt+\iint_{\mathscr{C}}V_{\textnormal{r}}p(\bm{x})\bar{\bm{v}}^{\mathrm{T}}(\bm{x},s)\mathbf{Q}_{2}(\bm{x},s)\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\hskip 17.00024pt+\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S),\end{split}$ (97)
@@
$\mathbf{Q}_{2}(\bm{x},s)$ $\triangleq-\mathbf{H}\bigl(\bar{\bm{v}}(\bm{x},s),s\bigr)^{\mathrm{T}}\mathbf{\Sigma}_{1}=\mathbf{M}^{2}\bigl(V_{\textnormal{r}}\bar{\bm{v}}(\bm{x},s)\bigr)\mathbf{G}^{-1}\bigl(V_{\textnormal{r}}\bar{\bm{v}}(\bm{x},s)\bigr)\mathbf{\Sigma}_{1}.$ (98b)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)+\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (99)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)-\dfrac{1}{2}\iint_{\mathscr{C}}\bigl(\nabla_{\bm{x}}\cdot p(\bm{x})\bar{\bm{V}}(\bm{x})\bigr)\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (100)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\dot{W}\bigl(\bm{z}(\cdot,s)\bigr),\qquad s\in(0,S).\end{split}$ (101)
Integrating the above Eq. (101) proves (70) with $w(\bm{f}(\cdot,s),-\bar{\bm{v}}(\cdot,s))=-\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}_{\textnormal{r}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}=-\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}$. ∎